OpenStudy (anonymous):

find the equation of the tangent line to f(x)=x^3 where x=2

OpenStudy (anonymous):

ok, we'll need the derivate of the function f'(x)=3x^2 now we can calculate the slope at x=2 f'(2)=3*2^2=12 now we need the point at x=2 f(x)=2^3=8 now a line looks like this: y = ax + b we have a (the slope) =12 we have x = 2 we have y =8 now we substitue and find b 8 = 12*2 +b b = -16 so the equation of the line is y = 12x - 16